# Visuospatial navigation without distance, prediction, integration, or maps

### v3

@@ Abstract
Navigation is controlled by at least two partially dissociable, concurrently developed systems in the brain. The cognitive map informs an organism of its location and bearing, updated by distance-based prediction and vestibular integration. Response-based systems, on the other hand, directly evaluate movement decisions from immediate percepts. Here we demonstrate the sufficiency of visual response-based decision-making in a classic open field navigation task often assumed to require a cognitive map. Three distinct strategies emerge to robustly navigate to a hidden goal, each conferring contextual tradeoffs, as well as aligning with behavior observed with rodents, insects, fish, and sperm cells. We propose reframing navigation from the bottom-up, without assuming online access to computationally expensive top-down representations, to better explain behavior under energetic or attentional constraints.
@@ 1 Introduction
Navigation involves correlating perceivable spatial features with expected goal location, whether it be food, shelter, or other area of interest, requiring the environment to be sensed and processed before effective movement. Recently published models for spatial navigation tend to employ either predictive or integrative architecture for building cognitive maps, which can then used for planning trajectories to desired goals [1, 2, 3, 4, 5]. Although these maps can be used for flexible planning, particularly regarding shortcuts, focus on this subset of navigational possibility has left its alternative relatively understudied.
The striatal response-based navigation system, partially dissociable from and concurrently learned alongside the hippocampal map-based system at individually variable preferential ratios [6, 7, 8, 9], can provide trusted routes to known destinations at lower latency, allowing navigational reliability while attention and energy can be directed to more critical or less predictable tasks [10, 11]. In addition to the striatum, the posterior parietal and retrosplenial cortices may also help compose egocentric movement choices along routes [12, 13].
@@
Map-based navigation algorithms often require online visual calculation and prediction of egocentric distances [3, 4, 5], or otherwise vestibular integration [1, 2], in order to function. Although visual distance calculation may be highly useful in specific contexts, e.g. cliff avoidance, prey capture, or jumping gaps [14], whether in the natural environment or in designed experimental paradigms, this perceptual dimension may not be as universal as assumed, or as accessible without considering its cost. Rats, for example, prioritize overhead binocular fusion over constant frontal depth perception, as the latter is not as critical as predator avoidance [15]. Furthermore complex 3D habitats may be required to justify its development [16]. In a similar vein, the primate tendency to fixate our eyes rather than heads during self-motion may prioritize visual conflict detection over vestibular integration [17, 18, 19].
@@
From an intuitive perspective, on our way to work or the market, routes we know quite well, to what extent do we continually update distances to landmarks, or count our turns, rather than rely on relative positions of cues in our visual field to make turn decisions? That allocentric map-based abilities both take longer to acquire and subside faster than egocentric navigation over the span of a lifetime suggests the preeminence of response-based decision-making in the context of constrained cognitive capacity [20, 21]. Rather than assuming unabated access to higher level representations and abilities, perhaps a simpler feedforward framework can explain a subset of navigational behaviors across the biological spectrum, which, given nature’s propensity to minimize energy, may be favored if performance is adequately robust.
@@
In pursuit of this alternative hypothesis, minimal visual perception-action loops of embodied agents, without predictive feedback control, are trained with fixed environment, perceptual, cognitive, and action constraints to solve a classic spatial navigation task. By minimizing assumptions specific to certain organisms, our conceptual approach seeks to maximize generalizability to discover fundamental properties of visual response-based navigation.
Feedforward response-based navigation is shown to be sufficient for open field navigation, a task often thought to require map-based knowledge [9, 22]. Two distinct behavioral classes emerge for minimal vision lacking distance information, quantitatively separable and preferentially developing at different rates with respect to visual resolution. A third class appears as a result of adding distance information into the visual input, an encoding approach describing greater perceptual or representational complexity, resulting in optimal trajectories despite lacking the common navigation representational assumptions.
@@ 2 Results
The environment is designed to be minimally complex: a square grid with four distinct walls and a circular foraging patch or goal. Using boundary walls restrains ability to estimate distance, forcing reliance on relative angles to grid corners as relevant landmarks. Agents are initialized in random locations and orientations for each simulation. As intermediate temporal representations are purposefully left out in our study, they must instead rely on immediate visuospatial calculations to robustly reach the patch.
@@
Figure 1: Agent perception-action loop flow (clockwise from bottom left): visual encoding, information processing, action conversion, environment update. Visual encoding consists of identifying walls corresponding to retinal angles of a raycast ($\upsilon$ number rays between -$\theta$ $\theta$ field of view limits) for minimal angle-only vision, then adding distance information according to scaling factor $\sigma$. Visual information passes through convolutional neural network, perceptron, linear output layer, and hyperbolic tangent tranformations to directly represent turning angle as well as speed via a linear function, which updates agent position and orientation for the next timestep (single update shown in environment and black point on linear function).
@@
Inspired by [23], we simulated visual input as a raycast matrix, one-hot encoded per ray, extending from the agent to the boundary walls (Fig. 1). Minimal vision provides information of identity and relative retinal angle of preceived walls, which we refer to as angle-based vision. For more complex vision, distance to the identified wall is logarithmically scaled into the encoding according to the Weber-Fechner law. In an unknown environment, accurate distance estimation requires higher order computation, e.g. stereopsis afforded by binocular fusion, static cues, dynamic temporal inference, or memory. Angle-based vision represents minimal environmental information, constrained either from an estimation reliability or a cognitive load, attentional-based perspective. Vision with distance information, scaled in at increasing signal variance ranges, characterizes increasing perceptual ability or cognitive bandwidth, which may include predictive or integrative inferential capacities.
@@
The information processing flow of an agent consists of three modules: a convolutional network (CNN), a single-layer perceptron, and a linear output layer. A CNN is chosen due to its conceptual similarity and quantified correlation to the visual cortex [24]. A simple single layer feedforward downstream network was found to confer sufficient performance, whereas more complex architectures were found to provide minor additional benefit (Fig. S1). Finally, a single network output represents a ratio between turning angle and speed, a minimal action space constraint describing the need to slow down in order to turn.
@@
The network is optimized via an evolutionary strategies (ES) algorithm. ES uses population-based black-box optimization without explicit gradient calculation, unlike reinforcement learning (RL) which optimizes a single agent by using backpropagation on a continuous differentiable objective function [25, 26].
@@
Fitness, or performance, was measured as the time taken to reach the patch plus the remaining distance. The first term is the primary driver for judging how well the agent navigates and the second guides initial learning behavior. The relative weighting of each was varied in initial experiments but found to not make a significant difference.
@@
Training the networks to this navigation task revealed the sufficiency of response-based navigation in an open field (Fig. S1). While only visual angle input is needed to adequately learn the task, performance and speed of convergence improves relative to the salience of an added distance signal. At maximal distance input ($\sigma=1$), performance approaches the theoretical lower bound of perfect trajectories straight to the patch. Attenuating its signal gradually reduces utility of the distance information, describing a titration between angle-based and distance-based perceptual encodings.
Figure 2: Three navigational classes, movement behavior & correlations. Top row: global movement behavior of three individual evolutionary runs or agents, with angle-based vision for A/B and with added distance scaling for C ($\sigma=1$), each with an 8 ray visual resolution ($\upsilon=8$). Solid lines: single agent trajectories from unique initial positions and orientations, temporally colored from purple to yellow, density and darkness reflect common routes. Dotted lines: 4 example individual trajectories. Black circle: patch location. Bottom row: movement correlations. Top left: spatial heading profile with respect to initial heading and distance from patch center, inset displays same 4 example initializations as dotted lines in above plots. Top right: polar histogram of relative orientation for timesteps when agent is over 100 grid units away from patch center, with frequency proportional to area and color of bin. Bottom left: temporal persistence of initial heading angle, marked by decorrelation time (50% threshold, red dashed line), sinusoidal shape reflecting correlated oscillations. Bottom right: directedness, an information theoretic measure calculated for each spatial bin in the environment, heatmap colored from low purple to high yellow. See Fig. S2, S3 for other individuals.
On the behavioral side, a convergence into distinct classes was observed in the navigation trajectory space (Fig. 2 top row, movies S1-S3). Two separate angle-only navigational strategies emerged despite identical constraints, best described by the terms "indirect sequential" and "biased diffusive". Adding distance scaling to the visual input resulted in a third class called "direct pathing". Hybrid strategies exist, indicating a continuum among clusters (Fig. S8), where exact boundaries depend on classification criteria details, discussed in Methods.
@@
The indirect sequential algorithm or class describes agents that learn to travel indirect trajectories to the patch, composed of a sequence of straight segments (Fig. 2a). Elliptical arcs sweeping across the arena define transition points, scaffolding the route, which the agent uses to shift direction through a single sharp turn. Despite lacking distance perception, the agents learn to use visuospatial invariances afforded by the environmental structure in order to reach the patch. However which invariances to use and how to compose them into an effective strategy varies between runs (Fig. S2).
@@
The route-determining elliptical manifolds are governed by a variety of environmental, perceptual, and movement constraints, coordinated to the patch location (Fig. S4, S5, see Methods). Mechanistically, the agent learns to tie major turn decisions to specific views with respect to two adjacent corners. As one of these views is rotated about the central axis of the respective wall, its decision threshold takes an elliptical form.
@@
The movement profiles of the indirect sequential agents stand in contrast to the second class of algorithms, which can be distinguished by looping trajectories (Fig. 2b). While overall progress is biased directly toward the patch, these agents regularly spin or turn, sensing and acting upon a wider set of observations. The trajectory maps render clouds of agent trajectories with little apparent correlation among each other, revealing no distinct manifolds or sequences, instead the process is diffusive.
@@
Curving trajectories dynamically sample the entire range of orientations, obscuring movement patterns. Though a vector field of its action output shows how learned gradients, scaffolded by the same elliptical manifolds, drive overall movement toward the patch (Fig. S6), a mechanism that does not explain indirect sequential movement whose straight paths rely on a static subsample of potential orientations. From a perturbative point of view, adding noise to the visual angle disrupts reliable indirect sequential navigation, while the biased diffusive agents remain robust (Fig. S7). In other words, the agents traverse perception-action attractors approximately centered on the patch, structured via the same invariances perceived by the indirect sequential agents, yet their movement pattern differentiates how they are used.
Figure 3: Classification & relative evolvability and fitnesses. A: Navigational algorithmic classes, separated by decorrelation time and directedness, top: colors correspond to distance scaling factor ($\sigma$) and visual resolution ($\upsilon$), bottom: colors correspond to algorithmic classes (IS: indirect sequential, BD: biased diffusion, DP: direct pathing). B: Angle-based strategy bifurcation, left: relative rates of either angle-based classes evolving under different visual resolutions, right: relative fitnesses, revealing optima at $\upsilon=8$ for IS and $\upsilon=16$ for BD. C: Distance-based phase transition, similar plots as in B, varying distance scaling factors and $\upsilon=8$, showing clear fitness difference and emergence of the DP class.
@@
While the behavioral bifurcation between the two angle-based navigation strategies is qualitatively apparent, robust separability calls for quantified measures. General movement behavior can be compared by relating orientation to initial heading as an agent approaches the patch (Fig. 2a/b, bottom row), demonstrating the tendency for straight, sparse or spinning, disperse profiles. However, while these plots provide useful intuition, extracting robust separation between the two classes for the entire range of agents has proven to be difficult, as can be seen in the extended figure (Fig. S3a/b, bottom row, top plots).
Instead of tracking relative orientation in space, its temporal counterpart provides a cleaner split (Fig. 2 bottom row, bottom left). Although other differences can be observed, the time needed for half of the agents to decorrelate from their initial heading consistently differentiates the two agent types. Biased diffusive agents quickly turn away, whereas indirect sequential agents follow their initial heading longer, until reaching an elliptical manifold.
Another way to track spatial correlation is via directedness, described as the orientational predictability of an agent at a given point in space. Average directedness is evidently lower for the biased diffusive agents (Fig. 2 bottom row, bottom right). The measure quantifies observations made for the top correlation plots of Fig. 2: indirect sequential agents tend to stick to few directions, while biased diffusive agents are more directionally indiscriminate. Yet for the former, directedness is locally low near elliptical manifolds and the patch, reflecting the wider range of potential directions at defined decision points, whereas there are no consistent spatial structures seen for the latter.
@@
Two clusters emerge in the phase space of decorrelation time and directedness, reflecting the qualitative separation observed above (Fig. 3a, S9 top left). How such clustering arises can be seen by plotting average fitness values for each agent or run (Fig. S9 top right). Runs close to cluster centers tend toward the greatest fitness, while those farther tend to drop. Such gradients suggest the two clusters are attractors in a fitness landscape, where optimization pulls toward the centers.
Given the asymmetric effect of noise to visual angle, it was expected that visual resolution would be critical in regulating preferential evolvability between the two classes. Indeed it is (Fig. 3b): lower visual resolution favors learning an indirect sequential strategy, while higher favors biased diffusive. Whereas the indirect sequential mechanism involves distinguishing only few major transitions, biased diffusive perception needs to construct an accurate gradient with respect to the entire range of potential visual observations, locations, and orientations. Therefore, visual resolution, a parameter dictating the degree of environmental complexity able to be perceived, affects relative occurence between strategies accordingly.
Relative fitness with respect to visual resolution shows a similar pattern (Fig. 3b). Each navigational strategy confers similar optimal global performance, explaining why the behavioral bifurcation exists in this parameter space. However, visual resolution needed for the biased diffusive agent to meet this level is twice that of the indirect sequential, thus requiring twice the number of convolutional operations per timestep, as well as twice the complexity in the visuospatial affordance space (Fig. S5 4th row).
@@
Adding distance information to the visual input gives rise to a third class of algorithms that can travel directly to the patch (Fig. 2c). Each agent differs in the degree and arrangement of how common routes collapse as they approach the patch, though variation is slim compared with the two angle-based classes. The uniform directness of their routes is apparent in their relative heading profiles, long decorrelation times, and high directedness throughout the environment, with the exception of the patch area.
Consequently, distance-based navigation establishes a third cluster center in the correlation space, (Fig. 3a). due to the lack of behavioral variability possible, whereas many different movement patterns can be classified under either angle-based class.
As the distance signal is attenuated, relative occurrence of direct pathing agents decreases until the other two classes fully occupy the resulting space (Fig. 3c left). Given that a sizable fitness difference between distance-based and angle-based strategies remains even with attenuated signal (Fig. 3c right), the abrupt disappearance of direct pathing behavior suggests the networks are no longer able to effectively use the input.

### v4

# Visuospatial navigation from the bottom-up: without vestibular integration, distance prediction, or maps

@@ Abstract
Navigation is believed to be controlled by at least two partially dissociable systems in the brain. The cognitive map informs an organism of its location and bearing, updated by integrating vestibular self-motion or predicting distances to landmarks. Route-based navigation, on the other hand, directly evaluate sequential movement decisions from immediate percepts. Here we demonstrate the sufficiency of visual route-based decision-making in a classic open field navigation task often assumed to require a cognitive map. Three distinct strategies emerge to robustly navigate to a hidden goal, each conferring contextual tradeoffs analyzed at both neural and behavioral scales, as well as qualitatively aligning with behavior observed across the biological spectrum. We propose reframing navigation from the bottom-up, through an egocentric episodic perspective without assuming online access to computationally expensive top-down representations, to better explain behavior under energetic or attentional constraints.
@@ 1 Introduction
Navigation involves correlating perceivable spatial features with expected goal location, whether it be food, shelter, or other area of interest, requiring the environment to be sensed and processed before effective movement. Cognitive maps, widely accepted as the ultimate way many organisms represent space, describe a system of hippocampal cells tuned to distinct spatial elements collectively composing a Euclidean graph upon which direct pathways and novel shortcuts can be calculated via simple vector transformation epsteinCognitiveMapHumans2017; bellmundNavigatingCognitionSpatial2018a. Although maps can be used for useful top-down planning, particularly regarding shortcuts, focus on this subset of navigational possibility has left its alternatives relatively understudied. Route-based navigation, on the other hand, concatenates sequences of egocentric movement decisions based on local observations, which can offer greater efficiency in complex environments peerStructuringKnowledgeCognitive2021; parra-barreroMapSpatialNavigation2023, or afford low demand reliability allowing energy to be directed to more critical or less predictable tasks prietoPlanningAutopilotAssociative2023. Following familiar routes may rely on the bottom-up application of episodic memory igloiLateralizedHumanHippocampal2010; goodroeComplexNatureHippocampalStriatal2018; wangEgocentricAllocentricRepresentations2020; kunzNeuralCodeEgocentric2021; mooreLinkingHippocampalMultiplexed2021; parra-barreroMapSpatialNavigation2023, a second key hippocampal function, while weaving together a network of routes enable construction of a non-Euclidean topology via state transitions warrenNonEuclideanNavigation2019; lanGoaldirectedNavigationHumans2025.
Map-based navigation algorithms often require online distance prediction recanatesiPredictiveLearningNetwork2021; kesslerHumanNavigationStrategies2024; stocklLocalPredictionlearningHighdimensional2024, or otherwise vestibular integration baninoVectorbasedNavigationUsing2018; cuevaEmergenceGridlikeRepresentations2018, in order to function. However, key visual dimensions were found to be functionally irrelevant for the two algorithms among those listed that performed ablation studies baninoVectorbasedNavigationUsing2018; recanatesiPredictiveLearningNetwork2021, suggesting that vision provides a redundant information stream.
In contrast, many animals have been shown to principally rely on visual cues for reliable navigation. Gerbils and hamsters tend to use configurations of visible landmarks when well-lit, conspicuous, and reliable, otherwise resorting to their sense of self-motion collettLandmarkLearningVisuospatial1986; jefferyNeurobiologySpatialBehaviour2003. Similarly, the primate tendency to fixate our eyes rather than heads during self-motion may prioritize visual conflict detection over vestibular integration SlownessSparsenessLead; garzorzVisualVestibularConflictDetection2017, which would explain stronger hippocampal cell correlations to head orientation and visual perspective than to spatial position rollsSpatialViewCells1997; ekstromCellularNetworksUnderlying2003; jacobsSenseDirectionHuman2010; maoSpatialModulationHippocampal2021; pizaPrimacyVisionShapes2024. Even rats, with normally strong positional hippocampal correlations, lose spatial selectivity in virtual reality despite maintaining directionality aghajanImpairedSpatialSelectivity2015; acharyaCausalInfluenceVisual2016 and flexible navigation mooreLinkingHippocampalMultiplexed2021. That is, restricting the vestibular sense forces a sequential route-based strategy grounded in vision.
@@
Although visual distance calculation may be highly useful in specific contexts, e.g. cliff avoidance, prey capture, or jumping gaps saleemInteractionsRodentVisual2023, whether in the natural environment or in designed experimental paradigms, this perceptual dimension may not be as universal as assumed, or as accessible without considering its cost. Eliminating depth does not significantly affect visual orientation capability in hamsters etienneNavigationSmallMammal1992. Rats prioritize overhead binocular fusion over constant frontal depth perception, as the latter is not as critical for predator avoidance wallaceRatsMaintainOverhead2013. Additionally, complex 3D habitats may be required to justify its development hagbiHeightsPlainsHow2022.
@@
On our way to work or the market, routes we know quite well, to what extent do we continually update distances to landmarks, or count our turns, rather than rely on relative positions of cues in our visual field? That allocentric map-based abilities both take longer to acquire and subside faster than egocentric navigation over the span of a lifetime suggests the preeminence of route-based decision-making in the context of constrained cognitive capacity colomboEgocentricAllocentricSpatial2017; tansanSpatialNavigationChildhood2022. Rather than assuming unabated access to high dimensional representations and abilities, perhaps a simpler feedforward framework can explain a subset of navigational behaviors across the biological spectrum, which, given nature’s propensity to minimize energy, may be favored if performance is adequately robust.
@@
In pursuit of this hypothesis, minimal visual perception-action loops of embodied agents, without predictive or integrative feedback, are trained with fixed environment, perceptual, cognitive, and action constraints to solve a classic spatial navigation task. Although the Morris Water Maze with random initialization is often described as testing map- or place-based navigation strategy involving vector calculation with respect to the hidden goal wolbersChallengesIdentifyingNeural2014; eichenbaumRoleHippocampusNavigation2017; vikbladhHippocampalContributionsModelBased2019; parra-barreroMapSpatialNavigation2023, we intend to demonstrate how this task may be simply solved by egocentric visual observations and sequential routes. Furthermore, by minimizing assumptions specific to certain organisms, our conceptual approach seeks to maximize generalizability to discover fundamental properties of visual route-based navigation.
@@ 2 Results
The environment is designed to be minimally complex: a square grid with four distinct walls and a hidden circular foraging patch or goal. Using boundary walls restrains ability to estimate distance, forcing reliance on relative angles to grid corners as salient landmarks. Agents are initialized in random locations and orientations for each simulation. As intermediate temporal representations are purposefully left out in our study, they must instead rely on immediate visuospatial calculations to robustly reach the patch.
@@
Figure 1: Agent perception-action loop flow. Clockwise from bottom left: visual encoding, information processing, action conversion, environment update. Visual encoding consists of identifying walls corresponding to retinal angles of a raycast ($\upsilon$ number rays between -$\theta$ & $\theta$ field of view limits) for minimal angle-only vision, then adding distance information according to scaling factor $\sigma$. Visual information passes through convolutional neural network, perceptron, linear output layer, and hyperbolic tangent tranformations to directly represent turning angle as well as speed via a linear function, which updates agent position and orientation for the next timestep (single update shown in environment and black point on linear function).
@@
Inspired by bastienModelCollectiveBehavior2020, we simulated visual input as a raycast matrix, one-hot encoded per ray, extending from the agent to the boundary walls (Fig 1). Minimal vision provides information of identity and relative retinal angle of preceived walls, which we refer to as angle-based vision. For more complex vision, distance to the identified wall is logarithmically scaled into the encoding according to the Weber-Fechner law. In an unknown environment, accurate distance estimation requires higher order computation, e.g. stereopsis afforded by binocular fusion, static cues, dynamic temporal inference, or memory. Angle-based vision, lacking distance information, represents minimal environmental information, constrained either from an estimation reliability or a cognitive load, attentional-based perspective. Vision with distance information, scaled in at increasing signal variance ranges, characterizes increasing perceptual ability or cognitive bandwidth, which may include predictive or integrative inferential capacities.
@@
The information processing flow of an agent consists of three modules: a convolutional network (CNN), a single-layer perceptron, and a linear output layer. A CNN is chosen mainly for its image processing ability, though in part to its conceptual similarity and quantified correlation to the visual cortex lindsayConvolutionalNeuralNetworks2021. A simple single layer feedforward downstream network was found to confer sufficient performance, whereas more complex architectures were found to provide minor additional benefit (S1 Fig). Finally, a single network output represents a ratio between turning angle and speed, a minimal action space constraint describing the need to slow down in order to turn. Note, this network is intended to be only phenomenologically related to sensorimotor information flow, rather than for specific biological detail.
@@
The network is optimized via an evolutionary strategies (ES) algorithm. ES uses population-based black-box optimization without explicit gradient calculation, unlike reinforcement learning (RL) which optimizes a single agent by using backpropagation on a continuous differentiable objective function salimansEvolutionStrategiesScalable2017; majidDeepReinforcementLearning2021. Though both being mechanisms for optimization through perception-action-based interaction in the environment, they both result in the construction of useful learned representations anggrainiNeuralSignaturesReinforcement2018; bermudez-contrerasNeuroscienceSpatialNavigation2020; vijayabaskaranNavigationTaskAction2022.
@@
Fitness, or performance, was measured as the time taken to reach the patch plus the remaining distance if the patch was not reached by the end of the simulation. The first term is the primary driver for judging how well the agent navigates and the second guides initial learning behavior. The relative weighting of each was varied in initial experiments but found to not make a significant difference.
@@
Training the networks to this navigation task revealed the sufficiency of route-based navigation in an open field (S1 Fig). Although only visual angle input is needed to adequately learn the task, performance and speed of convergence improves relative to the salience of an added distance signal. At maximal distance input ($\sigma=1$), performance approaches the theoretical lower bound of perfect trajectories straight to the patch. Attenuating its signal gradually reduces utility of the distance information, describing a titration between angle-based and distance-based perceptual encodings. At $\sigma=0.2$ signal variation is too subtle for the network to effectively use the information, slowing down learning rate and hindering performance, while at $\sigma=0.1$ the signal is not evidently used at all.
Figure 2: Three navigational classes, movement behavior & correlations. Top row: global movement behavior of three individual evolutionary runs or agents, with angle-based vision for A/B and with added distance scaling for C ($\sigma=1$), each with an 8 ray visual resolution ($\upsilon=8$). Solid lines: single agent trajectories from unique initial positions and orientations, temporally colored, density and darkness reflect common routes. Dotted lines: 4 example individual trajectories. Black circle: patch location. Bottom row: movement correlations. Top left: spatial heading profile with respect to initial heading and distance from patch center, inset shows same 4 example trajectories in above plots. Top right: polar histogram of relative orientation for timesteps when agent is over 100 grid units away from patch center, with frequency proportional to area and color of bin. Bottom left: temporal persistence of initial heading angle, marked by decorrelation time (50% threshold, red dashed line), sinusoidal shape reflecting correlated oscillations. Bottom right: directedness heatmap, an information theoretic measure calculated for each spatial bin in the environment. See S2 and S3 Figs for other individuals.
On the behavioral side, a convergence into distinct classes was observed in the navigation trajectory space (Fig 2 top row, movies S1-S3). Two separate angle-only navigational strategies emerged despite identical constraints, best described by the terms "indirect sequential" and "biased diffusive". Adding distance scaling to the visual input resulted in a third class called "direct pathing". Hybrid strategies exist, indicating a continuum among clusters (S8 Fig), where exact boundaries depend on classification criteria details, discussed in Materials and methods.
@@
The indirect sequential algorithm or class describes agents that learn to travel indirect trajectories to the patch, composed of a sequence of straight segments (Fig 2a). Elliptical arcs sweeping across the arena define transition points, scaffolding the route, which the agent uses to shift direction through a single sharp turn. Despite lacking distance perception, the agents learn to use visuospatial invariances afforded by the environmental structure in order to reach the patch. However which invariances to use and how to compose them into an effective strategy varies between runs (S2 Fig).
@@
The route-determining elliptical manifolds are governed by a variety of environmental, perceptual, and movement constraints, coordinated to the patch location (S4 and S5 Figs, see Materials and methods). Mechanistically, the agent learns to tie major turn decisions to specific views with respect to two adjacent corners. As one of these views is rotated about the central axis of the respective wall, its decision threshold takes an elliptical form (S5 Fig, 1b). While these manifolds cannot be easily seen by examining static neural activity normalized across location and orientation (Fig 3A), the ellipses are clearly evident when using goal-directed temporal data (Fig 3D). The goal location does not appear to spatially correlate with uniform grid data yet strongly overlaps with the visible manifolds in trajectory data (Fig 3A & D, left). However, directional components are not evident when deriving neural tuning curves for the former while strong 90 degree components arise in the latter (Fig 3A & D, right).
@@
The movement profiles of the indirect sequential agents stand in contrast to the second class of algorithms, which can be distinguished by looping trajectories (Fig 2b). While overall progress is biased directly toward the patch, these agents regularly spin or turn, sensing and acting upon a wider set of observations. The trajectory maps render clouds of agent trajectories with little apparent correlation among each other, revealing no distinct manifolds or sequences, instead the process is diffusive.
@@
Curving trajectories dynamically sample the entire range of orientations, obscuring movement patterns. Though its neural representation (Fig 3B) and vector field of action output (S6 Fig) show how learned gradients, scaffolded by the same elliptical manifolds, drive overall movement toward the patch, a mechanism that does not explain indirect sequential movement whose straight paths rely on a limited subsample of potential orientations. Goal location does correlate with averaged grid data for these agents, while the trajectory data seems less useful to describe mechanism. From a perturbative point of view, adding noise to the visual angle disrupts reliable indirect sequential navigation, while the biased diffusive agents remain robust (S7 Fig). Taken together, the agents traverse perception-action attractors approximately centered on the patch, structured via the same invariances perceived by the indirect sequential agents, yet their movement pattern differentiates how they are used.
Figure 3: Distinct approaches to represent neural data. Top: neural activity data gathered with uniform spatial and orientational occupancy. Bottom: agents initialized as above, data gathered over 500 timesteps and binned in an equivalent manner. Left: spatial selectivity regardless of orientation, normalized from zero to maximum activation. Middle eight: spatial selectivity with respect to orientation, illustrated by agent directions. Right: neural tuning curves with respect to egocentric angle from the patch. A-C & D-F relate to the three agents and navigational classes of Fig 2: indirect sequential (IS), biased diffusive (BD), and direct pathing (DP).
@@
While the behavioral bifurcation between the two angle-based navigation strategies is qualitatively apparent, robust separability calls for quantified measures. General movement behavior can be compared by relating orientation to initial heading as an agent approaches the patch (Fig 2a/b, bottom row), demonstrating the tendency for straight, sparse or spinning, disperse profiles. However, though these plots provide useful intuition, extracting robust separation between the two classes for the entire range of agents has proven to be difficult, as can be seen in the extended figure (S3a/b Fig, bottom row, top plots).
Instead of tracking relative orientation in space, its temporal counterpart provides a cleaner split (Fig 2 bottom row, bottom left). Although other differences can be observed, the time needed for half of the agents to decorrelate from their initial heading consistently differentiates the two agent types. Biased diffusive agents quickly turn away, whereas indirect sequential agents follow their initial heading longer, until reaching an elliptical manifold.
Another way to track spatial correlation is via directedness, described as the orientational predictability of an agent at a given point in space. Average directedness is evidently lower for the biased diffusive agents (Fig 2 bottom row, bottom right). The measure quantifies observations made for the top correlation plots of Fig 2: indirect sequential agents tend to stick to few directions, while biased diffusive agents are more directionally indiscriminate. Yet for the former, directedness is locally low near elliptical manifolds and the patch, reflecting the wider range of potential directions at defined decision points, whereas there are no consistent spatial structures seen for the latter.
Figure 4: Classification & relative evolvability and fitnesses. A: Navigational algorithmic classes, separated by decorrelation time and directedness, top: colors correspond to distance scaling factor ($\sigma$) and visual resolution ($\upsilon$), bottom: colors correspond to algorithmic classes (IS: indirect sequential, BD: biased diffusion, DP: direct pathing), including hybrids as combined half-circles. B: Angle-based strategy bifurcation, left: relative rates of either angle-based classes evolving under different visual resolutions, right: relative fitnesses, revealing optima at $\upsilon=8$ for IS and $\upsilon=16$ for BD. C: Distance-based phase transition, similar plots as in B, varying distance scaling factors and $\upsilon=8$. Scatter plots on right include violins for sample size greater than 5.
@@
Two clusters emerge in the phase space of decorrelation time and directedness, reflecting the qualitative separation observed above (Fig 4a, S9 top left). How such clustering arises can be seen by plotting average fitness values for each agent or run (S9 Fig, top right). Runs close to cluster centers tend toward the greatest fitness, while those farther tend to drop. Such gradients suggest the two clusters are attractors in a fitness landscape, where optimization pulls toward the centers.
Nonlinear dimensionality reduction techniques can find the gradient between the two navigational classes (S10 Fig) directly via the set of all possible visual observations and actions in the environment, however with lower explainability than using the two correlation metrics.
Given the asymmetric effect of noise to visual angle, it was expected that visual resolution would be critical in regulating preferential evolvability between the two classes. Indeed it is (Fig 4b): lower visual resolution favors learning an indirect sequential strategy, while higher favors biased diffusive. Whereas the indirect sequential mechanism involves distinguishing only few major transitions, biased diffusive perception needs to construct an accurate gradient with respect to the entire range of potential visual observations, locations, and orientations. Therefore, visual resolution, a parameter dictating the degree of environmental complexity able to be perceived, affects relative occurence between strategies accordingly.
Relative fitness with respect to visual resolution shows a similar pattern (Fig 4b). Each navigational strategy confers similar optimal global performance, explaining why the behavioral bifurcation exists in this parameter space. However, visual resolution needed for the biased diffusive agent to meet this level is twice that of the indirect sequential, thus requiring twice the number of convolutional operations per timestep, as well as twice the complexity in the visuospatial affordance space (S5 Fig, 4th row).
@@
Adding distance information to the visual input gives rise to a third class of algorithms that can travel directly to the patch (Fig 2c). Each agent differs in the degree and arrangement of how common routes collapse as they approach the patch, though variation is slim compared with the two angle-based classes. The uniform directness of their routes is apparent in their relative heading profiles, long decorrelation times, and high directedness throughout the environment, with the exception of the patch area. Appropriately, their neural representations appear to outline clear directional trajectories to the patch position (Fig 3C/F) with oriented activations cleanly situated between environment boundaries and the patch. The goal position is outlined using either dataset, with strong orientational alignment when using grid data, appearing to represent a goal-direction cell sarelVectorialRepresentationSpatial2017 as well as behaviorally affording goal vectors regardless of initialization (Fig 3F, middle).
Distance-based navigation establishes a third cluster center in the correlation space, (Fig 4a). due to the lack of behavioral variability possible, whereas many different movement patterns can be classified under either angle-based class. As the distance signal is attenuated, relative occurrence of direct pathing agents decreases until the other two classes fully occupy the resulting space (Fig 4c left). Given that a sizable fitness difference between distance-based and angle-based strategies remains even with attenuated signal (Fig 4c right), the abrupt disappearance of direct pathing behavior suggests the networks are no longer able to effectively use the input.